# LEO-VL: Towards 3D Vision-Language Generalists via Data Scaling with Efficient Representation

### v1

@@ 5 Conclusion
We propose LEO-VL, an advanced 3D-VL generalist that excels at general 3D-VL tasks with natural language outputs. LEO-VL is equipped with an efficient scene representation that bridges 2D perception and 3D spatial structure, featuring strong perception capability, high representation efficiency, and simplified 3D spatial modeling. For the generalist learning of LEO-VL, we curate a comprehensive 3D-VL dataset that spans four domains of real-world indoor scenes (i.e., ScanNet, 3RScan, MultiScan, and ARKitScenes) and five tasks, covering captioning, QA, planning, and dialogue. We demonstrate the state-of-the-art performance of LEO-VL on various 3D QA benchmarks such as SQA3D. Ablative analyses reveal the importance of diversity in tasks and scene domains. The exploration with additional simplistic QA data confirms our principle that prioritizes data quality over scale, which naturally leads to consistent scaling effects. We hope our experience and findings in this work can benefit future research in developing 3D-VL generalists. Furthermore, we demonstrate the necessity and advantage of post-training objectives in enhancing the robustness of 3D-VL models.
@@ Appendix A Discussion
This work introduces a 3D-VL generalist model for 3D scene understanding tasks with natural language outputs, offering a step forward in enabling intelligent systems to comprehend and interact with the physical world. The proposed approach holds promise for applications such as assistive robotics, AR/VR interfaces, and digital twin systems for smart environments. Meanwhile, we recognize potential concerns related to data privacy, computational resource consumption, and misuse in surveillance. We encourage the community to pair technical innovation with ethical foresight. And we will release our code and data to support transparent, reproducible, and responsible research in 3D-VL generalists.
@@ Appendix B Implementation Details
We finetune the LLM using LoRA [30] for parameter-efficient learning, with a rank of 16, scaling factor , and a dropout rate of 0, following LEO [34]. LoRA adapters are applied to all projection matrices: in the attention layers and in the MLP blocks. For inference, we employ beam search with repetition and length penalties. Full hyperparameter settings are provided in Table 7.
@@
We present the detailed settings in Table 7. The instruction tuning of LEO-VL is conducted for 5 epochs, and the post-training is conducted for 10 epochs.
@@
We calculate GPT-Score using the Azure OpenAI API with GPT-4o-2024-08-06 as the evaluator. Each score is evaluated based on a triplet consisting of a question, a ground-truth answer, and a predicted answer. We adopt the scoring prompt from MSR3D [46], which includes a system instruction and example contexts to guide the evaluation criteria.
@@ Appendix C Additional Results
Scan2Cap [14] is a benchmark for 3D object captioning, which utilizes the referential text from ScanRefer [9] as captions. We have not included Scan2Cap results in the main paper as the captions are limited in linguistic diversity and offer limited insight into a model’s descriptive capability. Nonetheless, for completeness, we report the performance of LEO-VL finetuned on Scan2Cap in Appendix C. The results show that LEO-VL achieves state-of-the-art performance on Scan2Cap, attaining the highest scores on two out of four evaluation metrics, highlighting its strong capability in object captioning.
@@
Scan2Cap [14] | 35.2 | 22.4 | 21.4 | 43.5 |
3DJCG [6] | 49.5 | 31.0 | 24.2 | 50.8 |
3D-VisTA [117] | 66.9 | 34.0 | 27.1 | 54.3 |
Vote2Cap-DETR++ [12] | 74.4 | 37.2 | 26.2 | 53.3 |
3DVLP [103] | 54.4 | 34.1 | 34.3 | 54.3 |
LEO [34] | 72.4 | 38.2 | 27.9 | 58.1 |
LL3DA [11] | 65.2 | 36.8 | 26.0 | 55.1 |
PQ3D [118] | 80.3 | 36.0 | 29.1 | 57.9 |
Chat-Scene [32] | 77.2 | 36.3 | 28.0 | 58.1 |
LLaVA-3D [115] | 79.2 | 41.1 | 30.2 | 63.4 |
3D-LLaVA [21] | 78.8 | 36.9 | 27.1 | 57.7 |
Video-3D LLM [111] | 80.0 | 40.2 | 28.5 | 61.7 |
GPT4Scene [71] | 86.3 | 40.6 | 28.2 | 59.3 |
@@
We present extended results with more metrics for the experiments reported in the main paper. We report the full results for ScanQA in Appendix C and SQA3D in Appendix C.
@@
MSQA.
We extend the baseline coverage on the MSQA benchmark. Existing baselines for MSQA are relatively scarce, primarily because models generally require training on MSQA to attain reasonable performance, i.e., zero-shot results tend to be poor. As only a handful of recent models have been trained on MSQA, the available baselines are limited. In Appendix C, we report additional baselines and present more detailed metrics on the MSQA ScanNet subset, including GPT-4o [60], LEO [34], MSR3D [46], and SplatTalk [83]. Note that existing evaluation results on MSQA include two different versions of test data: the initially released version and a subsequent refined version. Results corresponding to the earlier version are marked with †.
Model | Count. | Exist. | Attr. | Spatial | Navi. | Others | Overall |
GPT-4o∗‡ [60] | 32.3 | 79.3 | 79.0 | 37.0 | 31.7 | 91.6 | 52.3 |
LEO [34] | 32.5 | 88.5 | 58.7 | 44.2 | 39.6 | 81.4 | 54.8 |
MSR3D [46] | 32.3 | 93.1 | 50.0 | 46.5 | 54.1 | 75.6 | 54.2 |
SplatTalk†‡ [83] | 19.6 | 60.3 | 44.0 | 35.8 | 35.5 | 61.8 | 41.8 |
LEO-VL (ScanNet) | 32.0 | 92.7 | 59.2 | 57.3 | 48.1 | 83.5 | 58.4 |
LEO-VL | 36.3 | 93.8 | 59.2 | 56.0 | 44.6 | 83.5 | 58.2 |
@@
Similar to MSQA, the Beacon3D benchmark also has evaluation results based on two versions of test data: an initial version corresponding to the results reported in the paper [35], and a refined version that is released. We denote results corresponding to the earlier version with †. In Appendix C, we report the results on Beacon3D with more detailed metrics.
@@
3D-VisTA† [117] | 20.5 | 33.5 | 52.1 | 33.8 | 36.5 | 35.3 | 8.1 |
PQ3D [118] | 37.8 | 45.8 | 32.1 | 19.2 | 44.5 | 35.9 | 4.2 |
SceneVerse [39] | 26.4 | 40.4 | 40.0 | 35.0 | 54.1 | 40.5 | 4.7 |
LEO [34] | 16.4 | 39.8 | 47.6 | 52.8 | 54.3 | 45.2 | 7.5 |
Chat-Scene† [32] | 36.4 | 39.8 | 56.7 | 47.6 | 48.8 | 45.8 | 7.8 |
@@ Introduction.
We adopt RoPE to encode the voxel heights in 3D scenes considering its advantage in capturing height information during voxel feature pooling. As illustrated in Fig. 9, when two objects swap their heights, RoPE-encoded features can distinguish the configurations, whereas additive position embeddings fail. This property motivates our adoption of RoPE for the encoding of height information.
@@
Fourier features [80, 44] focus on encoding multi-dimensional spatial positions, such as pixel positions in an image. Classical methods typically process each dimension independently and then combine via concatenation. However, this falls short in capturing holistic spatial relations such as Euclidean distances, and struggles with generalizing to unbounded ranges of multi-dimensional coordinates. Fourier features overcome these issues by mapping the low-dimensional coordinates into a high-dimensional periodic function space, enabling better modeling of multi-dimensional positions.
@@ Appendix E Data Statistics
In the main paper, we explore the outcome of naively scaling with low-quality 3D-VL data. The additional QA data is sourced from LEO [34] (94k from 3RScan) and MMScan [54] (575k from ScanNet and 3RScan). This part of data exhibits a simplistic distribution with limited answer diversity. By categorizing answers into placeholder-based templates, we observe a significantly higher proportion of the top-15 frequent answer templates in the additional QA data (39.2% vs. 12.8%). The templates and their distribution are visualized in Figs. 10 and 11.
@@ Appendix F Qualitative Examples
We present qualitative results of LEO-VL performing various tasks in Fig. 12.

### v2

@@ 5 Conclusion
We propose LEO-VL, an advanced 3D VLM that excels at general 3D-VL tasks with natural language outputs. LEO-VL is equipped with an efficient scene representation that features high efficiency, strong perception capability, and simplified 3D spatial modeling. As the representation efficiency unlocks scalability of 3D-VL learning, we curate a comprehensive 3D-VL dataset that spans four real-world indoor domains (i.e., ScanNet, 3RScan, MultiScan, and ARKitScenes) and five tasks, covering captioning, QA, planning, and dialogue. LEO-VL achieves state-of-the-art performance on various 3D QA benchmarks such as SQA3D and Beacon3D. Our extensive ablation studies reveal the importance of diversity in tasks and scene domains, as well as our data curation principle for effective scaling of 3D-VL learning. Furthermore, we demonstrate the advantages of our SceneDPO for the post-training of 3D VLM compared to SFT and GRPO. We hope our experience and findings can benefit future research in 3D VLMs.
Ethics Statement
This work adheres to the ICLR Code of Ethics. All datasets used in our experiments will be publicly available, and no personally identifiable or sensitive data is involved. Synthetic data generated by LLM is used in a limited and controlled manner to augment training, without introducing privacy, fairness, or safety risks. We believe our work does not pose foreseeable ethical concerns or harmful societal impacts beyond those common to general research in 3D VLMs.
Reproducibility Statement
We have provided concrete information to ensure the reproducibility of our work. The model architecture, data scheme, and post-training objective are elaborated in Section˜3. Implementation configurations are presented in Sections˜4 and B. We will release the code and data upon acceptance to facilitate full reproducibility of our results.
@@ Appendix A Discussion
This work introduces a 3D VLM for scene understanding tasks, offering a step forward in enabling intelligent systems to comprehend and interact with the physical world. The proposed approach holds promise for applications such as assistive robotics, AR/VR interfaces, and digital twin systems for smart environments. Meanwhile, we recognize potential concerns related to data privacy, computational resource consumption, and misuse in surveillance. We encourage the community to pair technical innovation with ethical foresight. And we will release our code and data to support transparent, reproducible, and responsible research in 3D VLMs.
LLM Usage Statement.
We use LLM in two ways: (1) to aid with polishing the writing, including improving grammar and coherence of the manuscript; and (2) to generate a portion of training data for our model. The LLM does not contribute to any significant part of this work.
@@ Appendix B Implementation Details
We finetune the LLM using LoRA (Hu et al., 2022) for parameter-efficient learning, with a rank of 16, scaling factor , and a dropout rate of 0, following LEO (Huang et al., 2024b). LoRA adapters are applied to all projection matrices: in the attention layers and in the MLP blocks. For inference, we employ beam search with repetition and length penalties. Full hyperparameter settings are provided in Table˜8.
@@
We present the detailed settings in Table˜8. The instruction tuning and post-training of LEO-VL are both conducted for 5 epochs.
@@
We calculate GPT-Score using the Azure OpenAI API with GPT-4o-2024-08-06 as the evaluator. Each score is evaluated based on a triplet consisting of a question, a ground-truth answer, and a predicted answer. We adopt the scoring prompt from MSR3D (Linghu et al., 2024), which includes a system instruction and example contexts to guide the evaluation criteria.
@@ Appendix C Additional Results
Scan2Cap (Chen et al., 2021) is a benchmark for 3D object captioning, which utilizes the referential text from ScanRefer (Chen et al., 2020) as captions. We have not included Scan2Cap results in the main paper as the captions are limited in linguistic diversity and offer limited insight into a model’s descriptive capability. Nonetheless, for completeness, we report the performance of LEO-VL finetuned on Scan2Cap in Appendix˜C. The results show that LEO-VL achieves state-of-the-art performance on Scan2Cap, attaining the highest scores on two out of four evaluation metrics, highlighting its strong capability in object captioning.
@@
| Scan2Cap (Chen et al., 2021) | 35.2 | 22.4 | 21.4 | 43.5 |
| 3DJCG (Cai et al., 2022) | 49.5 | 31.0 | 24.2 | 50.8 |
| 3D-VisTA (Zhu et al., 2023) | 66.9 | 34.0 | 27.1 | 54.3 |
| Vote2Cap-DETR++ (Chen et al., 2024b) | 74.4 | 37.2 | 26.2 | 53.3 |
| 3DVLP (Zhang et al., 2024a) | 54.4 | 34.1 | 34.3 | 54.3 |
| LEO (Huang et al., 2024b) | 72.4 | 38.2 | 27.9 | 58.1 |
| LL3DA (Chen et al., 2024a) | 65.2 | 36.8 | 26.0 | 55.1 |
| PQ3D (Zhu et al., 2024b) | 80.3 | 36.0 | 29.1 | 57.9 |
| Chat-Scene (Huang et al., 2024a) | 77.2 | 36.3 | 28.0 | 58.1 |
| LLaVA-3D (Zhu et al., 2024a) | 79.2 | 41.1 | 30.2 | 63.4 |
| 3D-LLaVA (Deng et al., 2025) | 78.8 | 36.9 | 27.1 | 57.7 |
| Video-3D LLM (Zheng et al., 2025) | 80.0 | 40.2 | 28.5 | 61.7 |
| GPT4Scene (Qi et al., 2025) | 86.3 | 40.6 | 28.2 | 59.3 |
@@
We present extended results with more metrics for the experiments reported in the main paper. We report the full results for ScanQA in Appendix˜C and SQA3D in Appendix˜C.
@@
We present detailed results with more metrics in Appendix˜C.
@@
| 3D-VisTA (Zhu et al., 2023) | 28.4 | 35.7 | 41.6 | 48.0 | 55.0 | 43.2 | 7.3 |
| PQ3D (Zhu et al., 2024b) | 37.8 | 45.8 | 32.1 | 19.2 | 44.5 | 35.9 | 4.2 |
| SceneVerse (Jia et al., 2024) | 26.4 | 40.4 | 40.0 | 35.0 | 54.1 | 40.5 | 4.7 |
| LEO (Huang et al., 2024b) | 16.4 | 39.8 | 47.6 | 52.8 | 54.3 | 45.2 | 7.5 |
| Chat-Scene (Huang et al., 2024a) | 30.0 | 42.7 | 50.0 | 53.9 | 62.9 | 49.8 | 10.9 |
| LLaVA-3D (Zhu et al., 2024a) | 35.1 | 66.7 | 62.5 | 54.2 | 62.9 | 59.1 | 19.0 |
| Video-3D LLM (Zheng et al., 2025) | 40.1 | 64.1 | 60.6 | 55.3 | 64.1 | 59.0 | 17.9 |
| GPT4Scene (Qi et al., 2025) | 38.1 | 59.7 | 59.3 | 52.6 | 66.1 | 57.2 | 17.9 |
@@ Introduction.
We adopt RoPE to encode the voxel heights in 3D scenes considering its advantage in capturing height information during voxel feature pooling. As illustrated in Fig.˜9, when two objects swap their heights, RoPE-encoded features can distinguish the configurations, whereas additive position embeddings fail. This property motivates our adoption of RoPE for the encoding of height information.
@@
Fourier features (Tancik et al., 2020; Li et al., 2021) focus on encoding multi-dimensional spatial positions, such as pixel positions in an image. Classical methods typically process each dimension independently and then combine via concatenation. However, this falls short in capturing holistic spatial relations such as Euclidean distances, and struggles with generalizing to unbounded ranges of multi-dimensional coordinates. Fourier features overcome these issues by mapping the low-dimensional coordinates into a high-dimensional periodic function space, enabling better modeling of multi-dimensional positions.
@@ Appendix E Data Statistics
In the main paper, we explore the outcome of naively scaling with low-quality 3D-VL data. The additional QA data is sourced from LEO (Huang et al., 2024b) (94k from 3RScan) and MMScan (Lyu et al., 2024) (575k from ScanNet and 3RScan). This part of data exhibits a simplistic distribution with limited answer diversity. By categorizing answers into placeholder-based templates, we observe a significantly higher proportion of the top-15 frequent answer templates in the additional QA data (39.2% vs. 12.8%). The templates and their distribution are visualized in Figs.˜10 and 11.
@@ Appendix F Qualitative Examples
We present qualitative results of LEO-VL performing various tasks in Fig.˜12.